If one ball is selected at random, determine the odds in favor of it being a ball other than the 8 ball.
Answer by stanbon(57219)

If one ball is selected at random, determine the odds in favor of it being a ball other than the 8 ball.
----------------
There are 15 balls.
P(8-ball) = 1/15
P(other) = 14/15
---
odds in favos of "other" = P(other)/P(8-ball) = 14/1 = 14:1
===================
Cheers,
Stan H.